Grade 12th passMechanics

A ball thrown down the incline strikes a point on incline 25 m below the horizontal. If ball rises to a maximum height of 20m above the point of projection find the angle of projection with the horizontal

Kumar Sambhav
7 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
Dear student 
20=u^2sin^2x/2g⇒usinα=20
−25=20t−1/210t^2⇒t=5s
75=(ucosα)t
Solve to get tanα=4/3
good luck